# Reinforcement Learning with Action Sequence for Data-Efficient Robot Learning

### v1

@@ Abstract
Training reinforcement learning (RL) agents on robotic tasks typically requires a large number of training samples. This is because training data often consists of noisy trajectories, whether from exploration or human-collected demonstrations, making it difficult to learn value functions that understand the effect of taking each action. On the other hand, recent behavior-cloning (BC) approaches have shown that predicting a sequence of actions enables policies to effectively approximate noisy, multi-modal distributions of expert demonstrations. Can we use a similar idea for improving RL on robotic tasks? In this paper, we introduce a novel RL algorithm that learns a critic network that outputs Q-values over a sequence of actions. By explicitly training the value functions to learn the consequence of executing a series of current and future actions, our algorithm allows for learning useful value functions from noisy trajectories. We study our algorithm across various setups with sparse and dense rewards, and with or without demonstrations, spanning mobile bi-manual manipulation, whole-body control, and tabletop manipulation tasks from BiGym, HumanoidBench, and RLBench. We find that, by learning the critic network with action sequences, our algorithm outperforms various RL and BC baselines, in particular on challenging humanoid control tasks.
Project website: younggyo.me/cqn-as.
Figure 1: Summary of results. Coarse-to-fine Q-Network with Action Sequence (CQN-AS) is a value-based RL algorithm that learns a critic network with action sequence. We study CQN-AS on 53 robotic tasks from BiGym (Chernyadev et al., 2024), HumanoidBench (Sferrazza et al., 2024), and RLBench (James et al., 2020), where prior model-free RL algorithms struggle to achieve competitive performance. We show that CQN-AS outperforms various RL and BC baselines such as CQN (Seo et al., 2024), DrQ-v2+ (Yarats et al., 2022), SAC (Haarnoja et al., 2018), and ACT (Zhao et al., 2023).
@@ 1 Introduction
Reinforcement learning (RL) holds the promise of continually improving policies through online trial-and-error experiences (Sutton & Barto, 2018), making it an ideal choice for developing robots that can adapt to various environments. However, despite this promise, training RL agents on robotic tasks typically requires a prohibitively large number of training samples (Kalashnikov et al., 2018; Herzog et al., 2023), which becomes problematic as deploying robots often incurs a huge cost. Therefore many of the recent successful approaches on robot learning have been based on behavior-cloning (BC; Pomerleau 1988), which can learn strong policies from offline expert demonstrations (Brohan et al., 2023b; a; Zhao et al., 2023; Chi et al., 2023; Team et al., 2024; Fu et al., 2024a).
One cause for the poor data-efficiency of RL algorithms on robotic tasks is that training data consists of noisy trajectories. When collecting data for training RL agents, we typically inject some noise into actions for exploration (Sehnke et al., 2010; Lillicrap et al., 2016) that may induce trajectories with jerky motions. Moreover, we often initialize training with human-collected demonstrations that can consist of noisy multi-modal trajectories (Chernyadev et al., 2024). Such noisy data distributions make it difficult to learn value functions that should understand the consequence of taking each action. We indeed find that prior RL algorithms perform much worse than the BC baseline on mobile bi-manual manipulation tasks with human-collected demonstrations when compared to a tabletop manipulation setup with synthetic demonstrations collected via motion-planning (see Figure 1).
@@
On the other hand, recent BC approaches have shown that predicting a sequence of actions enables policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Inspired by this, in this paper, we investigate how to use a similar idea for improving the data-efficiency of RL algorithms on robotic tasks. In particular, we present a novel RL algorithm that learns a critic network that outputs Q-values over a sequence of actions (see Figure 2). By training the critic network to explicitly learn the consequence of taking a series of current and future actions, our algorithm enables the RL agents to effectively learn useful value functions from noisy trajectories. We build this algorithm upon a recent value-based RL algorithm that learns RL agents to zoom-into continuous action space in a coarse-to-fine manner (Seo et al., 2024), thus we refer to our algorithm as Coarse-to-fine Q-Network with Action Sequence (CQN-AS).
@@
To evaluate the generality and capabilities of CQN-AS, we study CQN-AS on various setups with sparse and dense rewards, and with or without demonstrations. In BiGym benchmark (Chernyadev et al., 2024), which provides human-collected demonstrations for mobile bi-manual manipulation tasks, CQN-AS outperforms various model-free RL and BC baselines (Yarats et al., 2022; Zhao et al., 2023; Seo et al., 2024). Moreover, in HumanoidBench (Sferrazza et al., 2024), which consists of densely-rewarded humanoid control tasks, we show that CQN-AS can also be effective without demonstrations, outperforming prior model-free RL baselines (Haarnoja et al., 2018; Seo et al., 2024). Finally, in RLBench (James et al., 2020), which provides synthetic demonstrations generated via motion-planning, CQN-AS achieves similar performance as model-free RL and BC baselines on most tasks, but significantly better performance on several long-horizon manipulation tasks.
Figure 2: Coarse-to-fine Q-network with action sequence. (Left) Our key idea is to train a critic network to output Q-values over a sequence of actions. We design our architecture to first obtain features for each sequence step and aggregate features from multiple sequence steps with a recurrent network. We then project these outputs into Q-values at level $l$. (Right) For action inference, we repeat the procedure of computing Q-values for level $l\in\{1,...,L\}$. We then find the action sequence with the highest Q-values from the last level $L$, and use it for controlling robots at each time step.
2 Background
@@
We mainly consider a robotic control problem which we formulate as a partially observable Markov decision process (Kaelbling et al., 1998; Sutton & Barto, 2018). At each time step $t$, an RL agent encounters an observation $\mathbf{o}_{t}$, executes an action $\mathbf{a}_{t}$, receives a reward $r_{t+1}$, and encounters a new observation $\mathbf{o}_{t+1}$ from the environment. Because the observation $\mathbf{o}_{t}$ does not contain full information about the internal state of the environment, in this work, we use a stack of past observations as inputs to the RL agent by following the common practice in Mnih et al. (2015). For simplicity, we omit the notation for these stacked observations. When the environment is fully observable, we simply use $\mathbf{o}_{t}$ as inputs. Our goal in this work is to train a policy $\pi$ that maximizes the expected sum of rewards through RL while using as few samples as possible, optionally with access to a modest amount of expert demonstrations collected either by motion-planners or by humans.
@@
Coarse-to-fine Q-Network (CQN; Seo et al. 2024) is a value-based RL algorithm for continuous control that trains RL agents to zoom-into the continuous action space in a coarse-to-fine manner. In particular, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient value-based RL algorithms (Mnih et al., 2015; Silver et al., 2017; Schrittwieser et al., 2020), designed to be used with discrete actions, for continuous control.
Formally, let $\mathbf{a}_{t}^{l}$ be an action at level $l$ with $\mathbf{a}_{t}^{0}$ being the zero vector [^1] For simplicity, we describe CQN and CQN-AS with a single-dimensional action in the main section. See Appendix B for full description with $N$-dimensional actions, which is straightforward but requires more indices.. We then define the coarse-to-fine critic to consist of multiple Q-networks which compute Q-values for actions at each level $\mathbf{a}_{t}^{l}$, given the features $\mathbf{h}_{t}$ and actions from the previous level $\mathbf{a}_{t}^{l-1}$, as follows:
$Q_{\theta}^{l}(\mathbf{h}_{t},\mathbf{a}_{t}^{l},\mathbf{a}_{t}^{l-1})\;\;\text{for}\;\;l\in\{1,...,L\}$ (1)
We optimize each Q-network at level $l$ with the following objective:
$\mathcal{L}^{l}=\left(Q_{\theta}^{l}(\mathbf{h}_{t},\mathbf{a}_{t}^{l},\mathbf{a}_{t}^{l-1})-r_{t+1}-\gamma\max_{a^{\prime}}Q_{\bar{\theta}}^{l}(\mathbf{h}_{t+1},a^{\prime},\pi^{l}(\mathbf{h}_{t+1})\right),$ (2)
where $\bar{\theta}$ are delayed parameters for a target network (Polyak & Juditsky, 1992) and $\pi^{l}$ is a policy that outputs the action $\mathbf{a}_{t}^{l}$ at each level $l$ via the inference steps with our critic, i.e., $\pi^{l}(\mathbf{h}_{t})=\mathbf{a}_{t}^{l}$. Specifically, to output actions at time step $t$ with the critic, CQN first initializes constants $a_{t}^{\tt{low}}$ and $a_{t}^{\tt{high}}$ with $-1$ and $1$. Then the following steps are repeated for $l\in\{1,...,L\}$:
@@
We then use the last level’s action $\mathbf{a}_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix B.
3 Coarse-to-fine Q-Network with Action Sequence
We present Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that learns a critic network that outputs Q-values for a sequence of actions $\mathbf{a}_{t:t+K}=\{\mathbf{a}_{t},...,\mathbf{a}_{t+K-1}\}$ for a given observation $\mathbf{o}_{t}$. Our main motivation for this design comes from one of the key ideas in recent behavior-cloning (BC) approaches, i.e., predicting action sequences, which helps resolve ambiguity when approximating noisy, multi-modal distributions of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Similarly, by explicitly learning Q-values of both current and future actions from the given state, our approach aims to mitigate the challenge of learning Q-values with noisy trajectories from exploratory behaviors or human-collected demonstrations.
This section describes how we design our critic network with action sequence (see Section 3.1) and how we utilize action sequence outputs to control robots at each time step (see Section 3.2). The overview of our algorithm is available in Figure 2.
@@ 3.1 Coarse-to-fine Critic with Action Sequence
Our key idea is to design a critic network to explicitly learn Q-values for current action and future actions from the current time step $t$, i.e., $\{Q(\mathbf{o}_{t},\mathbf{a}_{t}),Q(\mathbf{o}_{t},\mathbf{a}_{t+1}),...,Q(\mathbf{o}_{t},\mathbf{a}_{t+K-1})\}$, to enable the critic to understand the consequence of executing a series of actions from the given state.
Formulation and objective
Let $\mathbf{a}_{t:t+K}^{l}=\{\mathbf{a}_{t}^{l},...,\mathbf{a}_{t+K-1}^{l}\}$ be an action sequence at level $l$ and $\mathbf{a}_{t:t+K}^{0}$ be a zero vector. We design our coarse-to-fine critic network to consist of multiple Q-networks that compute Q-values for each action at sequence step $k$ and level $l$:
$Q_{\theta}^{l,k}(\mathbf{h}_{t},\mathbf{a}_{t+k-1}^{l},\mathbf{a}_{t:t+K}^{l-1})\;\text{for}\;\;l\in\{1,...,L\}\;\text{and}\;k\in\{1,...,K\}$ (3)
We optimize our critic network with the following objective:
$\sum\nolimits_{k}\sum\nolimits_{l}\left(Q_{\theta}^{l,k}(\mathbf{h}_{t},\mathbf{a}_{t+k-1}^{l},\mathbf{a}_{t:t+K}^{l-1})-r_{t+1}-\gamma\max_{a^{\prime}}Q_{\bar{\theta}}^{l,k}(\mathbf{h}_{t+1},a^{\prime},\pi^{l}_{K}(\mathbf{h}_{t+1})\right)^{2},$ (4)
where $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$. In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible because Q-values for future actions depend only on current features $\mathbf{h}_{t}$ but not on Q-values for previous actions. We find this simple design, with independence across action sequence, works well even on challenging humanoid control tasks with high-dimensional action spaces (Sferrazza et al., 2024). We expect our idea can be strengthened by exploiting the sequential structure, i.e., Q-values at subsequent steps depend on previous Q-values (Metz et al., 2017; Chebotar et al., 2023), but we leave it as future work.
@@
We implement our critic network to initially extract features for each sequence step $k$ and aggregate features from multiple steps with a recurrent network (see Figure 2). This architecture is often helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},\mathbf{a}_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (Cho et al., 2014) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l}))$. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $Q_{\theta}^{l,k}(\mathbf{o}_{t},\mathbf{a}_{t+k-1}^{l},\mathbf{a}_{t:t+K}^{l-1})=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.

### v2

# Coarse-to-fine Q-Network with Action Sequence for Data-Efficient Robot Learning

@@ Abstract
In reinforcement learning (RL), we train a value function to understand the long-term consequence of executing a single action. However, the value of taking each action can be ambiguous in robotics as robot movements are typically the aggregate result of executing multiple small actions. Moreover, robotic training data often consists of noisy trajectories, in which each action is noisy but executing a series of actions results in a meaningful robot movement. This further makes it difficult for the value function to understand the effect of individual actions. To address this, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a novel value-based RL algorithm that learns a critic network that outputs Q-values over a sequence of actions, i.e., explicitly training the value function to learn the consequence of executing action sequences. We study our algorithm on 53 robotic tasks with sparse and dense rewards, as well as with and without demonstrations, from BiGym, HumanoidBench, and RLBench. We find that CQN-AS outperforms various baselines, in particular on humanoid control tasks.
Project webpage: younggyo.me/cqn-as
Pieter Abbeel
University of California, Berkeley
mail@younggyo.me
@@ 1 Introduction
Reinforcement learning (RL) holds the promise of continually improving policies through online experiences (Sutton & Barto, 2018). However, training RL agents on robotic tasks often requires a prohibitively large number of training samples (Kalashnikov et al., 2018; Herzog et al., 2023), which is problematic as deploying robots incurs a huge cost.
We posit that one cause for the poor data-efficiency of RL for robotics is the credit assignment problem (Minsky, 1961). In value-based RL, we typically train a value function to understand the long-term consequences of taking individual actions. However, this becomes challenging in robotics as robot movements are the aggregate result of a sequence of low-level actions, where the effect of each action is often ambiguous. For instance, moving a robot arm forward involves multiple steps of changing joint angles. While one may easily understand the consequence of executing a series of actions, it is often difficult to learn how each low-level action, such as changing the joint angle by certain degrees, contributes to the resulting movement.
Figure 1: Noisy trajectories in robotic data. We visualize the $(x,y,z)$ coordinates of a gripper in demonstrations from RLBench (James et al., 2020) that uses motion-planning for generating demonstrations and BiGym (Chernyadev et al., 2024) that provides human-collected demonstrations. (Right) We report the success rates of replaying BiGym demonstrations with various action smoothing schemes. We find that naïve approach of smoothing actions can make demonstrations be invalid as smoothed actions often lose precision, highlighting the need for developing RL algorithms that can learn from noisy robotic training data.
Moreover, robotic training data often consists of noisy trajectories, amplifying the credit assignment problem. For instance, we often initialize training with human-collected demonstrations that consist of noisy multi-modal trajectories (Chernyadev et al., 2024). In such trajectories, each action may be noisy because of human errors; however, executing a sequence of actions results in meaningful robot movement. One may think that this issue can be resolved by smoothing actions, but it often makes demonstrations be invalid by reducing action precision (see Figure 1). Moreover, when training RL agents, we typically inject some noise into actions for exploration (Sehnke et al., 2010; Lillicrap et al., 2016), which induces trajectories with jerky motions. This characteristic of robotic training data further makes it difficult to learn the value function that can properly evaluate the long-term consequence of each individual action.
On the other hand, recent behavior cloning (BC; Pomerleau 1988) approaches have shown that predicting a sequence of actions enables policies to effectively approximate the noisy, multi-modal distribution of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Inspired by this, we investigate whether we can similarly train a value function to explicitly learn the consequence of executing action sequences.
As a straightforward implementation of this idea, in Figure 2, we train actor-critic algorithms (Haarnoja et al., 2018; Fujimoto et al., 2018) with action sequence on stand task from HumanoidBench (Sferrazza et al., 2024). Specifically, we train the actor to output action sequence and the critic to take action sequence as inputs instead of single-step actions. Unfortunately, we observe that actor-critic algorithms with action sequence suffer from a severe value overestimation. This is because a wider action space makes the critic more vulnerable to function approximation error (Fujimoto et al., 2018) and the actor can easily exploit the estimation error.
@@
These results motivate us to design our RL algorithm with action sequence upon a recent value-based algorithm, i.e., Coarse-to-fine Q-Network (CQN; Seo et al. 2024), which solves continuous control tasks with discrete actions. We find that training this critic-only algorithm with action sequence enables us to achieve the benefits of RL with action sequence yet avoid the value overestimation problem. In particular, we introduce Coarse-to-fine Q-Network with Action Sequence (CQN-AS), which learns a critic network that outputs Q-values over a sequence of actions (see Figure 3). By training the critic network to explicitly learn the consequence of taking a sequence of current and future actions, CQN-AS enables the RL agents to effectively learn useful value functions on challenging robotic tasks.
@@
We show that CQN-AS outperforms various baselines on diverse setups with sparse and dense rewards, as well as with or without demonstrations: (i) mobile bi-manual manipulation tasks from BiGym (Chernyadev et al., 2024) that provides human-collected demonstrations, (ii) densely-rewarded humanoid control tasks from HumanoidBench (Sferrazza et al., 2024), and (iii) table-top manipulation tasks from RLBench (James et al., 2020) that provides synthetic demonsrations generated via motion-planning.
Figure 2: Training actor-critic algorithms with action sequences leads to value overestimation. We report target Q-values recorded throughout training actor-critic algorithms, i.e., SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018), with action sequence on $\tt{stand}$ task from HumanoidBench (Sferrazza et al., 2024).
2 Preliminaries
@@
We formulate a robotic control problem as a partially observable Markov decision process (Kaelbling et al., 1998; Sutton & Barto, 2018). At time step $t$, an RL agent encounters an observation $\mathbf{o}_{t}$, executes an action $a_{t}$, receives a reward $r_{t+1}$, and encounters a new observation $\mathbf{o}_{t+1}$ from the environment. We aim to train a policy $\pi$ that maximizes the expected sum of rewards through RL while using as few online samples as possible, optionally with access to a modest amount of expert demonstrations.
@@
(a) Coarse-to-fine inference procedure
(b) Architecture
Figure 3: Coarse-to-fine Q-Network with Action Sequence (CQN-AS). We build our algorithm upon Coarse-to-fine Q-Network (CQN; Seo et al. 2024), a recent critic-only RL algorithm that solves continuous control tasks with discrete actions. (a) In CQN framework, we train RL agents to zoom-into the continuous action space by iterating the procedures of (i) discretizing continuous action space into $B$ bins and (ii) find the bin with the highest Q-value to further discretize at the next level. We then use the last level’s action sequence for controlling robots. CQN-AS extends this idea to action sequences by computing actions for all sequence steps $k\in[1,...,K]$ in parallel. (b) We train a critic network to output Q-values over a sequence of actions. We design our architecture to first obtain features for each sequence step and aggregate features from multiple sequence steps with a recurrent network. We then project these outputs into Q-values.
@@
Coarse-to-fine Q-Network (CQN; Seo et al. 2024) is a value-based RL algorithm that solves continuous control tasks with discrete actions. The main idea of CQN is to train an RL agent to learn to select coarse discrete actions in shallower levels with larger bin sizes, and then refine their choices by selecting finer-grained actions in deeper levels with smaller bin sizes. Specifically, CQN iterates the procedures of (i) discretizing the continuous action space into multiple bins and (ii) selecting the bin with the highest Q-value to further discretize. This reformulates the continuous control problem as a multi-level discrete control problem, allowing for the use of ideas from sample-efficient discrete RL algorithms (Mnih et al., 2015; Silver et al., 2017) for continuous control.
Formally, let $a_{t}^{l}$ be an action at level $l$ with $a_{t}^{0}$ being the zero vector. [^1] For simplicity, we describe CQN and CQN-AS with a single-dimensional action in the main section. See Appendix B for full description with $N$-dimensional actions. We then define the coarse-to-fine critic to consist of multiple Q-networks which compute Q-values for actions at each level $a_{t}^{l}$, given the features $\mathbf{h}_{t}$ and actions from the previous level $a_{t}^{l-1}$, as follows:
$Q^{l}_{\theta}(\mathbf{h}_{t},a^{l-1}_{t})=\left[Q^{l}_{\theta}(\mathbf{h}_{t},a^{l,b}_{t},a^{l-1}_{t})\right]_{b=1}^{B}\in\mathbb{R}^{B}$
where $B$ is the number of bins for each level. We note that CQN uses scalar values representing the center of each bin for $a^{l-1}_{t}$, enabling the network to locate itself without access to all previous levels’ actions. We optimize each Q-network at level $l$ with the following objective:
[…]
where $\bar{\theta}$ are delayed parameters for a target network (Polyak & Juditsky, 1992) and $\pi^{l}$ is a policy that outputs the action $a_{t}^{l}$ at each level $l$ via the inference steps with our critic, i.e., $\pi^{l}(\mathbf{h}_{t})=a_{t}^{l}$. Specifically, to output actions at time step $t$, CQN first initializes constants $a_{t}^{\tt{low}}$ and $a_{t}^{\tt{high}}$ with $-1$ and $1$. Then the following steps are repeated for $l\in\{1,...,L\}$:
@@
We then use the last level’s action $a_{t}^{L}$ as the action at time step $t$. For more details, including the inference procedure for computing Q-values, we refer readers to Appendix B.
3 Method
We present Coarse-to-fine Q-Network with Action Sequence (CQN-AS), a value-based RL algorithm that learns a critic network that outputs Q-values for a sequence of actions $a_{t:t+K}=\{a_{t},...,a_{t+K-1}\}$ for a given observation $\mathbf{o}_{t}$. Our main motivation comes from one of the key ideas in recent BC approaches: predicting action sequences, which helps resolve ambiguity when approximating noisy distributions of expert demonstrations (Zhao et al., 2023; Chi et al., 2023). Similarly, by explicitly learning Q-values of a sequence of actions from the given state, our approach mitigates the challenge of learning Q-values with noisy trajectories. We provide the overview of CQN-AS in Figure 3.
Figure 4: Examples of robotic tasks. We study CQN-AS on 53 tasks spanning mobile bi-manual manipulation, whole-body control, and tabletop manipulation from BiGym (Chernyadev et al., 2024), HumanoidBench (Sferrazza et al., 2024), and RLBench (James et al., 2020).
@@ 3.1 Coarse-to-fine Critic with Action Sequence
Objective
Let $a_{t:t+K}^{l}=\{a_{t}^{l},...,a_{t+K-1}^{l}\}$ be an action sequence at level $l$ and $a_{t:t+K}^{0}$ be a zero vector. We design our coarse-to-fine critic network to consist of multiple Q-networks that compute Q-values for each action at sequence step $k\in\{1,...,K\}$ and level $l\in\{1,...,L\}$:
$Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t:t+K}^{l-1})=\left[Q_{\theta}^{l,k}(\mathbf{h}_{t},a_{t+k-1}^{l,b},a_{t:t+K}^{l-1})\right]_{b=1}^{B}$
where $B$ is the number of bins for each level. We optimize our critic network with the following objective:
[…]
where $N$ is a hyperparameter for $n$-step return and $\pi_{K}^{l}$ is an action sequence policy that outputs the action sequence $\mathbf{a}_{t:t+K}^{l}$ by following the similar inference procedure as in Section 2 (see Figure 3(a)). In practice, we compute Q-values for all sequence step $k\in\{1,...,K\}$ in parallel, which is possible as Q-values for future actions depend only on features $\mathbf{h}_{t}$ but not on previous actions.
Remarks on objective with $N$-step return
We note that any $N$-return can be used in Equation 1 because the network can learn the long-term value of outputting action $a_{t+k}$ from bootstrapping. There is a trade-off: if one considers a short $N$-step return, it can cause a challenge as the setup becomes a delayed reward setup; but training with higher $N$-step return may introduce variance (Sutton & Barto, 2018). In our considered setups, we empirically find that using common values $N\in\{1,3\}$ works the best. We provide empirical analysis on the effect of $N$ in Figure 8(c).
@@
Our critic network initially extracts features for each sequence step $k$ and aggregates features from multiple steps with a recurrent network (see Figure 3(b)). This architecture is helpful in cases where a single-step action is already high-dimensional so that concatenating them make inputs too high-dimensional. Specifically, let $\mathbf{e}_{k}$ denote an one-hot encoding for $k$. At each level $l$, we construct features for each sequence step $k$ as $\mathbf{h}^{l}_{t,k}=\left[\mathbf{h}_{t},a_{t+k-1}^{l-1},\mathbf{e}_{k}\right]$. We then encode each $\mathbf{h}_{t,k}^{l}$ with a shared MLP network and process them through GRU (Cho et al., 2014) to obtain $\mathbf{s}^{l}_{t,k}=f^{\tt{GRU}}_{\theta}(f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,1}^{l}),...,f^{\tt{MLP}}_{\theta}(\mathbf{h}_{t,k}^{l})$. We find that this design empirically performs better than directly giving actions as inputs to GRU. We then use a shared projection layer to map each $\mathbf{s}_{t,k}^{l}$ into Q-values at each sequence step $k$, i.e., $Q_{\theta}^{l,k}(\mathbf{o}_{t},a_{t+k-1}^{l},a_{t:t+K}^{l-1})=f_{\theta}^{\tt{proj}}(\mathbf{s}_{t,k}^{l})$.